A watermelon is launched from a 25-foot tall platform at a carnival. The watermelon's height h (in feet) at time t seconds can be modeled by the equation h = -16t^2 + 56t + 25. Find the maximum height of the watermelon

1
Expert's answer
2021-05-12T01:52:27-0400

We have given the equation of height,


"h(t) = -16t^2+56t+25"


"h'(t) = -32t+56"

For maximum height,


"h'(t) = 0"


Hence,


"-32t+56 = 0"


"t = \dfrac{56}{32} = 1.75s"


at "t = 1.75s"


Maximum Height "= -16(1.75)^2+56(1.75)+25"


"= -28+98+25"


"= 95feet"